OpenStudy (anonymous):

can someone label a b c d in this function pls Let y=−3sin[5(x−π/6)]+5

OpenStudy (anonymous):

y=asin[b(x-c)]+d a=Amplitude -3 negative so 'flips' curve b/2π = period = 5/2π radians to go through 1 full cycle c=shift in x-axis so sin term is 0 when x=π/6 d=shift in y-axis curve move up by 5